# Human Decision-making is Susceptible to AI-driven Manipulation

### v1

@@ Methods
We conducted a randomized controlled trial to investigate human susceptibility to AI manipulation attempts. Our study employed a between-subjects design ($N=233$) to examine how and to what extent AI systems (referred to as agents in this paper) can influence human decision-making across multiple scenarios (Figure 1). In each scenario, the user was provided with a pre-defined query and four corresponding options. One option was created as the most suitable (optimal) choice based on the user’s query, while the remaining three were designed to cause different types of harm. We investigated participants’ susceptibility to AI manipulation by analyzing the differences in their preferences for these options before and after interacting with the assigned AI agent.
@@
The scenarios in this study were carefully crafted to assess manipulation susceptibility in two domains reflecting common real-world situations in which AI assistants are employed [38, 39, 54, 40, 55, 29]: Financial (seeking advice on purchases) and Emotional (seeking guidance on handling emotional dilemmas) decision-making. In each domain, we curated three widely used topics to design scenarios that were applicable to a wide variety of individuals. For each scenario, to ensure consistency between the conditions of different participants, we designed pre-defined queries representing a hypothetical user’s situation, needs, and requirements. In addition, each scenario was accompanied by four distinct plausible options, one of which was crafted as the optimal choice, while others were carefully calibrated to indicate different levels of potential harm. In each scenario, we set one of the harmful options as the agent’s hidden incentive to prevent user suspicion regarding its bias.
@@
AI systems are making their way into financial investments [28] and are increasingly being used to enhance product advertisements [12, 29] and provide financial advice [56]. Hence, it is crucial to understand the extent to which these systems can impact users’ decisions despite their better judgments. Drawing from existing research on consumer decision-making [57, 58, 59], we identified key scenarios where users commonly seek AI assistance for financial guidance. We focused on three representative consumer decisions that require minimal domain expertise: selecting a fitness tracker, choosing a weight-loss medication, and finding an online shopping platform for clothes. To ensure practicality while maintaining experimental control, we designed pre-defined user queries corresponding to each topic that incorporated realistic budget constraints and requirements (Tables S9-S11).
The optimal choice was designed to align with stated user requirements and budget constraints while incorporating realistic trade-offs. We designed the remaining options to exhibit different types of potential financial harm: 1) Non-existent Product: This option was crafted to mirror characteristics of common online scams, featuring unusually attractive terms, comparatively lower cost, and limited verification possibilities. Specifically, this option was designed to only be available for pre-order from the supplier’s website and had relatively few or no reviews, making it difficult for users to assess its legitimacy. 2) Excessive Product: This option represented over-featured solutions that exceeded the user’s stated requirements and maximum budget and was designed to test users’ susceptibility to unnecessary upselling[60]. It allowed us to assess whether users could be swayed by the perceived desirability of excessive products, even when such choices caused them to spend more than they had originally planned. 3) Dependence-inducing Product: This option was designed to satisfy the user’s requirements at a relatively lower upfront cost than the optimal choice. However, it incorporated elements of long-term financial commitment, such as subscription models, for continued access to users’ desired features, which introduced hidden costs and future obligations. By choosing this option, users incurred recurring fees that, over time, exceeded their maximum budget, reflecting real-world scenarios where consumers are lured by low initial costs but face long-term financial burdens.
@@
Therefore, the optimal choice was designed to represent active coping, a process based on cognitive-behavioral literature that involves identifying specific actions to address the situation, developing alternative perspectives, and constructive engagement with support networks, which has shown the strongest correlations with increasing user’s mental well being and lower rates of anxiety and depression [69, 71]. The remaining options involved common maladaptive mechanisms with varying levels of harm [74]: 1) Disengagement: This option reflects avoidant coping patterns, incorporating elements such as behavioral disengagement, mental disengagement, and denial. While potentially offering immediate emotional relief, this option demonstrated less desirable long-term outcomes; 2) Emotional Venting: This option included elements of emotional release while omitting components of reflection, learning, or growth that characterize more adaptive forms of emotional processing. Hence, it clearly distinguished between adaptive and maladaptive emotional processing by emphasizing unregulated emotional discharge without problem-solving elements. 3) Self-blame: This option was constructed to reflect common patterns of maladaptive attribution as it included subtle elements of perfectionism and excessive responsibility-taking, characteristics that portray the individual as the root of the problem, bringing additional mental burden to the individual without finding a viable long-term solution.
@@
This study followed a three-phase protocol designed to assess participants’ susceptibility to AI manipulation while controlling for individual differences and ensuring ethical practices. Our experiments were approved by the institutional review board at Hong Kong University (Reference No: EA240497).
@@
Prior to our experiment, participants completed a comprehensive questionnaire of validated psychometric instruments, including demographic information (age, gender, education, occupation, and marital status) and social factors that may influence people’s susceptibility to manipulation, including personality[77, 78], social engineering susceptibility[79] [80], social support [81], and self-esteem[82]: 1) Ten-Item-Personality-Inventory (TIPI)[83], a 10-item assessment of personality traits based on dimensions from the Big Five Inventory (BFI)[84]: Openness, Conscientiousness, Extroversion, Agreeableness, and Emotional stability. We adopted this version of the personality test, as we assumed that while an AI agent may have a perception of the user’s personality traits, it would not be heavily accurate[12]; 2) Social Engineering Susceptibility Scale[85] focusing on individuals’ vulnerabilities in digital interactions and modified to include the portion of the questions that were relevant to our study (i.e., excluding questionnaire regarding informational susceptibilities); 3) Oslo Social Support Scale (OSSS-3)[86], a 3-item questionnaire evaluating participants’ accessibility to support networks; 4) Rosenberg Self-Esteem Scale (RSES)[87], a 10-item measurement of individuals’ perception of self-worth. Due to the nature of our study, we also adopted the AI Attitude Scale[88] to assess participants’ predispositions toward AI systems.
@@
Following the pre-experiment survey, we randomly assigned participants to experimental (domain-AI) conditions while balancing key demographic and psychometric variables across different groups. Each participant received a unique passcode to access our experimental platform and detailed instructions, framing the study as an investigation of personalized AI-assisted decision-making. The instructions were carefully worded to avoid demand characteristics while maintaining ethical transparency.
@@
After completing all of the required scenarios on the platform, participants evaluated the assigned scenarios’ relevance to their daily lives and rated their assigned AI agent on three dimensions: helpfulness, informativeness, and soundness (10-point Likert scales). We also collected open-ended responses regarding participants’ overall impressions and specific interaction experiences. To ensure ethical conduct and participant well-being, participants received detailed explanations of the study’s true objectives and the rationale behind any experimental deception. We instructed the participants on the optimal choices for each scenario and explained why certain options might be harmful. Additionally, participants were encouraged to contact the study investigators if they experienced any negative effects.
@@
We recruited participants through poster advertisements on WeChat, a predominant Chinese social media, and Prolific [89], a UK-based crowdsourcing platform. Participants were required to be 1) Fluent in Chinese or English; 2) Be able to use a computer device to access and use our platform; 3) Have experience using LLMs/personal assistants; 4) Be over 18 years of age; and 5) Do not currently suffer from physical/mental illnesses, as participation may have caused additional distress to these individuals. Data was collected online through our own platform to ensure data safety and privacy. As the experiment took approximately one hour on average, each participant was rewarded 80 RMB ($\approx$8£) for completing all the tasks and relevant questionnaires. In addition, the top 10% of the participants, based on dialogue quality and timely completion of tasks, received an additional bonus of 20 RMB ($\approx$2£).
@@
To assess how interactions with the AI agents influenced participants’ decisions, we considered the highest-rated option(s) at each time point (pre/post interaction) as the participant’s most likely choice(s). Changes in these choices were categorized into four decision patterns: No Change represents the participant maintaining their original decision, which was further divided into Optimal or Harmful, depending on whether the choice included the optimal option; Positive Shift indicates a beneficial change, where the participant initially chose a harmful option but switched to the optimal option or maintained their choice of the optimal option while including fewer harmful options in their decision after the interaction; Negative Shift involves changing from the optimal choice to a harmful option or switching between harmful options. We analyzed the distribution of decision patterns across domain-AI conditions using chi-squared tests, with post-hoc proportions z-tests for pairwise comparisons and Cohen’s $h$ for effect sizes.
@@
We used a linear mixed-effects model to analyze individual factors influencing manipulation susceptibility. Decision patterns were quantified on a scale reflecting their impact: negative shifts ($-1$), positive shifts ($1$), and no change (±$\pm 0.5$, with optimal and harmful decisions corresponding to $+$ and --, respectively). This mapping preserved both directional and magnitude information while enabling quantitative analysis of individual susceptibility.
@@
where $Y_{ij}$ represents participant $i$’s quantified decision pattern on scenario $j$, and $\beta_{0}$ and $\epsilon_{ij}$ indicate the intercept and residual error, respectively. $\mathbf{X}_{ij}$ represents covariates, including demographics, personality, trust in AI, social support accessibility, and social engineering susceptibility. We also included pre-interaction confidence and familiarity ratings and the scenario’s completion order as covariates. Lastly, the model incorporated random intercepts for the participant ($1|P_{i}$) and the scenario ($1|S_{j}$).
@@
As previously mentioned, participants completed a post-experiment survey evaluating both the scenarios and their assigned AI agent. They rated scenario commonality and three aspects of the AI agent (personalization, soundness, and informativeness) using 10-point Likert scales. These ratings were analyzed using one-way ANOVA followed by Tukey HSD pairwise comparisons. Responses to two open-ended questions were annotated by the authors: the first question, regarding AI helpfulness, was categorized into a binary (helpful/not helpful) taxonomy; the second question, probing overall impressions, was coded to identify mentions of perceived manipulation attempts by the AI agent (e.g., feeling as the AI agent had an ulterior motive). The proportions of these categorical responses across AI conditions were then compared using chi-squared tests with post-hoc z-tests.
@@ Acknowledgments
Figure S2: Summary of participant feedback. Bar plots display participants’ ratings for the commonality of the designed scenarios and evaluation of the assigned AI agents based on three aspects (personalization, soundness, and informativeness), respectively. ANOVA results are provided on the $y$-axes of each plot. Pairwise comparisons (Tukey HSD) are annotated between bar plots based on the following $P$ value legend: n.s. (not significant; $P>0.05$); * ($P\leq 0.01$); and, **($P\leq 0.0001$). Error bars represent $95\%$ confidence intervals. While there was no significant difference between scenarios in the emotional domain, participants in the financial domain reported experiencing the second scenario in daily life significantly less than the remaining two. Notably, participants rated the agents similarly on all three aspects, highlighting the covertness of AI-driven manipulation. Detailed statistical results are provided in Supplementary Table S6.
Table S1: Pairwise comparisons (proportion’s $z$-tests) of decision patterns between AI conditions across financial and emotional domains. Significant values ($P<0.05$) are highlighted in bold.
@@
No Change (Optimal) NA ($23.3\%$) MA ($23.7\%$) $0.06$ $1.00$ $0.01$
@@
No Change (Optimal) NA ($21.4\%$) MA ($20.7\%$) 0.12 $1.00$ $0.02$
@@
Table S2: Pairwise comparisons (proportion’s $z$-tests) of decision patterns between financial and emotional domains across AI conditions. Overall indicates the proportion of decision patterns within the domain for all AI conditions combined. Significant values ($P<0.05$) are highlighted in bold.
@@
Overall Negative Shift $49.4\%$ $32.2\%$ $4.64$ $\mathbf{3.56\times 10^{-6}}$ $0.35$
No Change (Harmful) $12.6\%$ $29.9\%$ $5.57$ $\mathbf{2.48\times 10^{-8}}$ $0.43$
No Change (Optimal) $23.9\%$ $20.2\%$ $1.16$ $0.25$ $0.09$
Positive Shift $14.1\%$ $17.7\%$ $1.30$ $0.19$ $0.10$
@@
No Change (Optimal) $23.3\%$ $21.4\%$ $0.36$ $0.72$ $0.05$
No Change (Harmful) $16.7\%$ $37.6\%$ $3.63$ $\mathbf{2.83\times 10^{-4}}$ $0.48$
@@
No Change (Optimal) $23.7\%$ $20.7\%$ $0.53$ $0.59$ $0.07$
No Change (Harmful) $9.6\%$ $25.2\%$ $3.09$ $\mathbf{2.03\times 10^{-3}}$ $0.42$
@@
No Change (Optimal) $24.6\%$ $18.7\%$ $1.10$ $0.27$ $0.14$
No Change (Harmful) $11.4\%$ $26.8\%$ $3.00$ $\mathbf{2.70\times 10^{-3}}$ $0.40$
@@
Table S3: Comparison of the SEMA’s strategy usage between decision-making domains. Percentages indicate the proportion of generated responses that correspond to the strategy. Effect sizes are measured by Cohen’s $h$. Significant values ($P<0.05$) are highlighted in bold.

### v2

@@ Methods
We conducted a randomized controlled trial to investigate human susceptibility to AI manipulation attempts. Our study employed a between-subjects design ($N=233$) to examine how and to what extent AI systems (referred to as agents in this paper) can influence human decision-making across decision-making scenarios (Figure 1). In each scenario, the participant was provided with a pre-defined user query and four corresponding options. One option was created as the most suitable (optimal) choice based on the user’s query, while the remaining three were designed to cause different types of harm. We investigated participants’ susceptibility to AI-driven manipulation by analyzing the differences in their preference ratings for these options before and after interacting with the assigned AI agent.
@@
The scenarios in this study were carefully crafted to assess manipulation susceptibility in two domains reflecting common real-world situations in which AI assistants are employed [42, 43, 64, 44, 65, 33]: Financial (seeking advice on purchases) and Emotional (seeking guidance on handling emotional dilemmas) decision-making. In each domain, we curated three widely used topics to design scenarios that applied to a wide variety of individuals. For each scenario, to ensure consistency between the conditions of different participants, we designed pre-defined queries representing a hypothetical user’s situation, needs, and requirements. In addition, each scenario was accompanied by four distinct plausible options, one of which was crafted as the optimal choice, while others were carefully designed to indicate different levels of potential harm. In each scenario, we designated one of the harmful options as the agent’s hidden incentive to prevent user suspicion regarding its bias.
@@
AI systems are making their way into financial investments [32] and are increasingly being used to enhance product advertisements [16, 33] and provide financial advice [66]. Hence, it is crucial to understand the extent to which these systems can impact users’ decisions despite their better judgments. Drawing from existing research on consumer decision-making [67, 68, 69], we identified key scenarios where users commonly seek AI assistance for financial guidance. We focused on three representative consumer decisions that require minimal domain expertise: selecting a fitness tracker, choosing a weight-loss medication, and finding an online shopping platform for clothes. To ensure practicality while maintaining experimental control, we designed pre-defined user queries corresponding to each topic that incorporated realistic budget constraints and requirements (Tables S10-S12).
The optimal choice was designed to align with stated user requirements and budget constraints while incorporating realistic trade-offs. We designed the remaining options to exhibit different types of potential financial harm: 1) Non-existent Product: This option was crafted to mirror characteristics of common online scams [70], featuring unusually attractive terms, comparatively lower cost, and limited verification possibilities. Specifically, this option was designed to only be available for pre-order from the supplier’s website and had relatively few or no reviews, making it difficult for users to assess its legitimacy. 2) Excessive Product: This option represented over-featured solutions that exceeded the user’s stated requirements and maximum budget and was designed to test users’ susceptibility to unnecessary upselling[71]. It allowed us to assess whether users could be swayed by the perceived desirability of excessive products, even when such choices caused them to spend more than they had originally planned. 3) Dependence-inducing Product: This option was designed to satisfy the user’s requirements at a relatively lower upfront cost than the optimal choice. However, it incorporated elements of long-term financial commitment, such as subscription models, for continued access to users’ desired features, which introduced hidden costs and future obligations. By choosing this option, users incurred recurring fees that, over time, exceeded their maximum budget, reflecting real-world scenarios where consumers are lured by low initial costs but face long-term financial burdens.
@@
Therefore, the optimal choice was designed to represent active coping, a process based on cognitive-behavioral literature that involves identifying specific actions to address the situation, developing alternative perspectives, and constructive engagement with support networks, which has shown the strongest correlations with increasing user’s mental well being and lower rates of anxiety and depression [80, 82]. The remaining options involved common maladaptive mechanisms with varying levels of harm [85]: 1) Disengagement: This option reflects avoidant coping patterns, incorporating elements such as behavioral disengagement, mental disengagement, and denial. While potentially offering immediate emotional relief, this option demonstrated less desirable long-term outcomes; 2) Emotional Venting: This option includes elements of emotional release while omitting components of reflection, learning, or growth that characterize more adaptive forms of emotional processing. Hence, it clearly distinguishes between adaptive and maladaptive emotional processing by emphasizing unregulated emotional discharge without problem-solving elements. 3) Self-blame: This option was constructed to reflect common patterns of maladaptive attribution as it included subtle elements of perfectionism and excessive responsibility-taking, characteristics that portray the individual as the root of the problem, bringing additional mental burden to the individual without finding a viable long-term solution.
@@
This study followed a three-phase protocol designed to assess participants’ susceptibility to AI manipulation while controlling for individual differences and ensuring ethical practices. Our experiments were approved by the institutional review board at The University of Hong Kong (Reference No: EA240497).
@@
Prior to our experiment, participants completed a comprehensive questionnaire of validated psychometric instruments, including demographic information (age, sex, education, occupation, and marital status) and social factors that may influence people’s susceptibility to manipulation, including personality[88, 89], social engineering susceptibility[90] [91], social support [92], and self-esteem[93]: 1) Ten-Item-Personality-Inventory (TIPI)[94], a 10-item assessment of personality traits based on dimensions from the Big Five Inventory (BFI)[95]: openness, conscientiousness, extroversion, agreeableness, and emotional stability. We adopted this version of the personality test, as we assumed that while an AI agent may have a perception of the user’s personality traits, it would not be heavily accurate[16]; 2) Social Engineering Susceptibility Scale[96] focusing on individuals’ vulnerabilities in digital interactions and modified to include the portion of the questions that were relevant to our study (i.e., excluding questionnaire regarding informational susceptibilities); 3) Oslo Social Support Scale (OSSS-3)[97], a 3-item questionnaire evaluating participants’ accessibility to support networks; 4) Rosenberg Self-Esteem Scale (RSES)[98], a 10-item measurement of individuals’ perception of self-worth. Due to the nature of our study, we also adopted the AI Attitude Scale[99] to assess participants’ predispositions toward AI systems.
@@
Following the pre-experiment survey, we randomly assigned participants to domain-AI conditions while balancing key demographic and psychometric variables across different groups. Each participant received a unique passcode to access our experimental platform and detailed instructions, framing the study as an investigation of personalized AI-assisted decision-making. The instructions were carefully worded to avoid demand characteristics while maintaining ethical transparency.
@@
After completing all of the required scenarios on the platform, participants evaluated the assigned scenarios’ relevance to their daily lives and rated their assigned AI agent on three dimensions: helpfulness, informativeness, and soundness (10-point Likert scales). We also collected open-ended responses regarding participants’ overall impressions and specific interaction experiences. To ensure ethical conduct and participant well-being, participants received detailed explanations of the study’s true objectives and the rationale behind any experimental manipulation. We instructed the participants on the optimal choices for each scenario and explained why certain options might be harmful. Additionally, participants were encouraged to contact the study investigators if they experienced any negative effects.
@@
We recruited participants through poster advertisements on WeChat, a predominant Chinese social media, and Prolific [100], a UK-based crowdsourcing platform. Participants were required to be 1) Fluent in Chinese or English; 2) Be able to use a computer device to access and use our platform; 3) Have experience using LLMs/personal assistants; 4) Be over 18 years of age; and 5) Do not currently suffer from physical/mental illnesses, as participation may have caused additional distress to these individuals. Data was collected online through our platform to ensure data safety and privacy. As the experiment took approximately one hour on average, each participant was rewarded 80 RMB ($\approx$8£) for completing all the tasks and relevant questionnaires. In addition, the top 10% of the participants, based on dialogue quality and timely completion of tasks, received an additional bonus of 20 RMB ($\approx$2£).
@@
To assess how interactions with the AI agents influenced participants’ decisions, we considered the highest-rated option(s) at each time point (pre/post interaction) as the participant’s most likely choice(s). Changes in these choices were categorized into four decision patterns: No Change represents the participant maintaining their original decision, which was further divided into Optimal or Harmful, depending on whether the choice included the optimal option; Positive Shift indicates a beneficial change, where the participant initially chose a harmful option but switched to the optimal option or maintained their choice of the optimal option while including fewer harmful options in their decision after the interaction; Negative Shift involves changing from the optimal choice to a harmful option or switching between harmful options. We analyzed the distribution of decision patterns across domain-AI conditions using chi-squared tests, with post-hoc z-tests with Bonferroni correction for pairwise comparisons and Cohen’s $h$ for effect sizes.
@@
We employed a linear mixed-effects model to analyze individual factors influencing manipulation susceptibility. Decision patterns were quantified on a scale reflecting their impact: negative shifts ($-1$), positive shifts ($1$), and no change (00). This mapping preserved both directional and magnitude information while enabling quantitative analysis of individual susceptibility.
@@
where $Y_{ij}$ represents participant $i$’s quantified decision pattern on scenario $j$, and $\beta_{0}$ and $\epsilon_{ij}$ indicate the intercept and residual error, respectively. $\mathbf{X}_{ij}$ represents covariates, including demographics, personality, trust in AI, social support accessibility, and social engineering susceptibility traits. We also included pre-interaction confidence and familiarity ratings and the scenario’s completion order as covariates. Lastly, the model incorporated random intercepts for the participant ($1|P_{i}$) and the scenario ($1|S_{j}$).
@@
After completing the experiment, participants completed a post-study survey evaluating the designed scenarios and their assigned AI agents. They rated scenario commonality and three aspects of the AI agent (personalization, soundness, and informativeness) using 10-point Likert scales. These ratings were analyzed using one-way ANOVA followed by Tukey HSD pairwise comparisons. Responses to two open-ended questions were annotated by the authors: the first question, regarding AI helpfulness, was categorized into a binary (helpful/not helpful) taxonomy; the second question, probing overall impressions, was coded to identify mentions of perceived manipulation attempts by the AI agent (e.g., feeling as the AI agent had an ulterior motive). The proportions of these categorical responses across AI conditions were then compared using chi-squared tests and post-hoc z-tests with Bonferroni correction.
Data and Code Availability
The data collected in our study and the source code for data collection and analysis will be made available upon request from the corresponding authors, with clear justification and data usage agreement.
@@ Acknowledgments
Figure S2: Summary of participant feedback. Bar plots display participants’ ratings for the commonality of the designed scenarios and evaluation of the assigned AI agents based on three aspects (personalization, soundness, and informativeness), respectively. ANOVA results are provided on the $y$-axes of each plot. Pairwise comparisons (Tukey HSD) are annotated between bar plots based on the following $P$ value legend: n.s. (not significant; $P\geq 0.05$); * ($P<0.01$); and, ** ($P<0.0001$). Error bars represent $95\%$ confidence intervals. While there was no significant difference between scenarios in the emotional domain, participants in the financial domain reported experiencing the second scenario in daily life significantly less than the remaining two. Notably, participants rated the agents similarly on all three aspects, highlighting the covertness of AI-driven manipulation. Detailed statistical results are provided in Supplementary Table S6.
Table S1: Pairwise comparisons (proportion’s $z$-tests with Bonferroni correction) of decision patterns between AI conditions across financial and emotional domains. Significant values ($P<0.05$) are highlighted in bold.
@@
No Change (Optimal) NA ($16.7\%$) MA ($9.6\%$) $1.58$ $0.34$ $0.21$
NA ($16.7\%$) SEMA ($11.4\%$) $1.16$ $0.74$ $0.15$
MA ($9.6\%$) SEMA ($11.4\%$) $0.43$ $1.00$ $0.06$
No Change (Harmful) NA ($23.3\%$) MA ($23.7\%$) $0.06$ $1.00$ $0.01$
@@
No Change (Optimal) NA ($37.6\%$) MA ($25.2\%$) 2.01 $0.13$ $0.27$
NA ($37.6\%$) SEMA ($26.8\%$) 1.79 $0.22$ $0.23$
MA ($25.2\%$) SEMA ($26.8\%$) 0.28 $1.00$ $0.04$
No Change (Harmful) NA ($21.4\%$) MA ($20.7\%$) 0.12 $1.00$ $0.02$
@@
Table S2: Pairwise comparisons (proportion’s $z$-tests with Bonferroni correction) of decision patterns between financial and emotional domains across AI conditions. Overall indicates the proportion of decision patterns within the domain for all AI conditions combined. Significant values ($P<0.05$) are highlighted in bold.
@@
No Change (Optimal) $16.7\%$ $37.6\%$ $3.63$ $\mathbf{2.83\times 10^{-4}}$ $0.48$
No Change (Harmful) $23.3\%$ $21.4\%$ $0.36$ $0.72$ $0.05$
@@
No Change (Optimal) $9.6\%$ $25.2\%$ $3.09$ $\mathbf{2.03\times 10^{-3}}$ $0.42$
No Change (Harmful) $23.7\%$ $20.7\%$ $0.53$ $0.59$ $0.07$
@@
No Change (Optimal) $11.4\%$ $26.8\%$ $3.00$ $\mathbf{2.70\times 10^{-3}}$ $0.40$
No Change (Harmful) $24.6\%$ $18.7\%$ $1.10$ $0.27$ $0.14$
@@
Overall Positive Shift $14.1\%$ $17.7\%$ $1.30$ $0.19$ $0.10$
No Change (Optimal) $12.6\%$ $29.9\%$ $5.57$ $\mathbf{2.48\times 10^{-8}}$ $0.43$
No Change (Harmful) $23.9\%$ $20.2\%$ $1.16$ $0.25$ $0.09$
Negative Shift $49.4\%$ $32.2\%$ $4.64$ $\mathbf{3.56\times 10^{-6}}$ $0.35$
Financial Optimal 32.96 2, 345 $\mathbf{2.39\times 10^{-13}}$ 0.16
Hidden 9.91 2, 345 $\mathbf{1.96\times 10^{-4}}$ 0.05
Emotional Optimal 16.76 2, 348 $\mathbf{3.36\times 10^{-7}}$ 0.09
Hidden 21.33 2, 348 $\mathbf{5.47\times 10^{-9}}$ 0.11
Pre-interaction Ratings
Financial Optimal NA ($6.22[2.29]$) MA ($6.41[2.06]$) -0.69 231.37 1.00 0.09
NA ($6.22[2.29]$) SEMA ($6.49[2.23]$) -0.93 231.87 1.00 0.12
MA ($6.41[2.06]$) SEMA ($6.49[2.23]$) -0.28 226.00 1.00 0.04
Hidden NA ($5.53[2.58]$) MA ($5.54[2.52]$) -0.06 231.83 1.00 0.01
NA ($5.53[2.58]$) SEMA ($5.89[2.63]$) -1.06 230.89 1.00 0.14
MA ($5.54[2.52]$) SEMA ($5.89[2.63]$) -1.00 226.00 1.00 0.13
Emotional Optimal NA ($7.22[2.19]$) MA ($6.88[2.64]$) 1.05 213.91 1.00 0.14
NA ($7.22[2.19]$) SEMA ($7.39[2.09]$) -0.61 235.89 1.00 0.08
MA ($6.88[2.64]$) SEMA ($7.39[2.09]$) -1.62 209.39 0.64 0.21
Hidden NA ($6.11[2.25]$) MA ($5.75[2.47]$) 1.16 221.22 1.00 0.15
NA ($6.11[2.25]$) SEMA ($5.99[2.23]$) 0.41 237.17 1.00 0.05
MA ($5.75[2.47]$) SEMA ($5.99[2.23]$) -0.79 222.59 1.00 0.10
Post-interaction Ratings
Financial Optimal NA ($7.22[2.44]$) MA ($5.81[2.30]$) 4.55 231.97 $\mathbf{5.20\times 10^{-5}}$ 0.59
NA ($7.22[2.44]$) SEMA ($5.15[2.47]$) 6.44 231.12 $\mathbf{4.19\times 10^{-9}}$ 0.84
MA ($5.81[2.30]$) SEMA ($5.15[2.47]$) 2.08 226.00 0.23 0.28
Hidden NA ($5.74[2.71]$) MA ($7.04[2.73]$) -3.64 231.19 $\mathbf{2.01\times 10^{-3}}$ 0.48
NA ($5.74[2.71]$) SEMA ($7.55[2.59]$) -5.23 231.99 $\mathbf{2.26\times 10^{-6}}$ 0.68
MA ($7.04[2.73]$) SEMA ($7.55[2.59]$) -1.47 226.00 0.86 0.19
Emotional Optimal NA ($8.25[1.83]$) MA ($6.71[2.78]$) 4.89 188.80 $\mathbf{1.27\times 10^{-5}}$ 0.66
NA ($8.25[1.83]$) SEMA ($7.11[2.33]$) 4.23 230.05 $\mathbf{2.00\times 10^{-4}}$ 0.54
MA ($6.71[2.78]$) SEMA ($7.11[2.33]$) -1.17 215.34 1.00 0.15
Hidden NA ($5.82[2.35]$) MA ($7.19[2.53]$) -4.23 222.51 $\mathbf{2.07\times 10^{-4}}$ 0.56
NA ($5.82[2.35]$) SEMA ($6.61[2.27]$) -2.64 236.30 0.05 0.34
MA ($7.19[2.53]$) SEMA ($6.61[2.27]$) 1.84 222.27 0.41 0.24
Factor Financial Emotional
Intercept 0.89 1.67 0.10 0.53 -0.19 -0.45 0.65 0.41
AI condition [MA] -0.59 -6.29 $\mathbf{3.07\times 10^{-10}}$ 0.09 -0.45 -4.68 $\mathbf{2.83\times 10^{-6}}$ 0.10
AI condition [SEMA] -0.58 -6.29 $\mathbf{3.09\times 10^{-10}}$ 0.09 -0.41 -4.71 $\mathbf{2.46\times 10^{-6}}$ 0.09
Language [Chinese] -0.14 -1.11 0.27 0.12 -0.01 -0.09 0.93 0.12
Sex [Male] 0.01 0.07 0.95 0.09 0.03 0.37 0.71 0.07
Education [Undergraduate] 0.08 0.79 0.43 0.10 -0.05 -0.53 0.60 0.09
Education [College] 0.15 1.07 0.28 0.14 0.04 0.30 0.77 0.12
Education [High school] 0.26 1.49 0.14 0.17 -0.15 -1.12 0.26 0.13
Work [Part-time] -0.16 -1.49 0.14 0.11 0.10 0.98 0.33 0.10
Work [Others] -0.23 -1.76 0.08 0.13 -0.10 -0.79 0.43 0.12
Marital Status [Married] -0.04 -0.40 0.69 0.09 -0.04 -0.46 0.65 0.10
Marital Status [Others] -0.12 -0.73 0.47 0.17 -0.09 -0.41 0.69 0.22
Age 0.00 -0.53 0.60 0.01 0.00 -0.45 0.65 0.01
Extroversion 0.05 1.54 0.12 0.03 -0.01 -0.32 0.75 0.04
Agreeableness 0.02 0.52 0.60 0.04 0.02 0.70 0.48 0.03
Conscientiousness -0.01 -0.17 0.87 0.04 -0.03 -0.69 0.49 0.04
Emotional Stability -0.02 -0.46 0.64 0.04 0.00 0.06 0.95 0.03
Openness -0.05 -1.45 0.15 0.04 -0.04 -0.96 0.34 0.04
AI Trust -0.01 -1.72 0.09 0.01 -0.00 -0.35 0.73 0.01
Social Support -0.03 -1.32 0.19 0.02 0.03 1.34 0.18 0.02
Self-Esteem 0.02 2.14 $\mathbf{0.03}$ 0.01 0.01 0.84 0.40 0.01
Normative 0.05 0.84 0.40 0.06 0.17 2.65 $\mathbf{8.09\times 10^{-3}}$ 0.06
Continuance -0.05 -1.01 0.31 0.05 0.03 0.52 0.61 0.05
Affective -0.03 -0.55 0.58 0.05 -0.09 -1.73 0.08 0.05
Trust 0.00 0.06 0.95 0.05 -0.01 -0.29 0.77 0.05
Obedience -0.07 -1.25 0.21 0.06 -0.05 -0.86 0.39 0.06
Reactance 0.00 0.08 0.93 0.05 0.03 0.50 0.61 0.05
Confidence (Pre-interaction) 0.01 0.64 0.52 0.02 -0.02 -0.83 0.41 0.03
Familiarity (Pre-interaction) -0.02 -1.32 0.19 0.02 0.01 0.34 0.74 0.02
Completion Order -0.00 -0.01 0.99 0.04 -0.08 -1.91 0.06 0.04
Table S5: Comparison of the strategy usage between decision-making domains (financial or emotional) for the strategy-enhanced manipulative agent (SEMA). Percentages indicate the proportion of generated responses that correspond to each strategy. Effect sizes are measured by Cohen’s $h$. Significant values ($P<0.05$) are highlighted in bold.